A water faucet can fill a 10.0 gal bath tub in 6.70 minutes. Given: 1 gal = 231 in3. How long will it take to fill a hot tub that has a cavity dimension of 2.00-m x 3.00-m x 1.00 m with this faucet?

Need this in hours
Answer #1

Volume V = 10 gal = 231 in3

1 gal = 0.87443 m3

V = 10 x 0.87443 = 8.74433 gal 0.00 3.78541.im volume= 10 gal 000378541 m m3 Time (t = 6.70 min = 6.70x6o. t=.402 s flow rate Q 00037834 402 Q 9.4 164t, = 637186 329 Seconds ti 1061697 min 1061097 ti hours 60 to 17669 hows hours m3

V = 0.0037854118 m3

time t = 6.70 min = 6.70 x 60 = 402 seconds

Flowrate Q = V/t = 0.0000094164 m3/s

Cavity volume V1 = 2 x 3 x 1 = 6 m3

time taken t1 = V1 / Q = 6 / 0.0000094164 = 637186.1858 seconds = 10619 minutes =